# IRIF, Université Paris Cité, France

### v1

@@ 5.1 Evaluation Metrics
MNIST Weights Recovered $100\%$ $100\%$ $100\%$ $100\%$ $90.63\%$ $93.75\%$ $90.63\%$ $95.31\%$ $100\%$ $99.72\%$
$-1784\text{-}8^{(8)}\text{-}$ 1 (II) Time $3$m24242424s $7$m5555s $7$m37373737s $10$m41414141s $11$m46464646s $28$m29292929s $20$m38383838s $36$m46464646s $0.01$s -
Queries $2^{21.59}$ $2^{24.21}$ $2^{24.22}$ $2^{24.24}$ $2^{24.26}$ $2^{25.37}$ $2^{24.98}$ $2^{25.63}$ 00 -
[…]
MNIST Weights Recovered $100\%$ $100\%$ $100\%$ $99.61\%$ $99.22\%$ $97.27\%$ $98.83\%$ $95.31\%$ $100\%$ $99.83\%$
784-$-1784\text{-}16^{(8)}\text{-}$ 1 (III) Time $6$m17171717s $26$m13131313s $28$m57575757s $39$m33333333s $1$h39393939m $2$h13131313m $1$h48484848m $3$h56565656m $0.01$s -
Queries $2^{21.60}$ $2^{24.22}$ $2^{24.24}$ $2^{24.67}$ $2^{25.65}$ $2^{26.04}$ $2^{25.79}$ $2^{26.95}$ 00 -
Foerster et al. [13]
MNIST 784-$-1784\text{-}16^{(8)}\text{-}$ 1 Layer coverage $100\%$ $100\%$ $100\%$
Weights Recovered $100\%$ $100\%$ $\geq 37.50\%^{\dagger}$ $\geq 0\%^{\dagger}$ $100\%$
Time $2$h46464646m $7$m50505050s - >36 hours - - - >36 hours $0.01$s -
Queries $2^{22.36}$ $2^{19.01}$ 00 -
5.2 Taxonomy of Unrecovered Weights
How does the weight recovery rate relate to coverage? To answer this question we need to understand more about the weights the attack does not recover. For example, as explained earlier (see Section 3.3), some weights are dead weights, meaning they can be set to zero without affecting the coverage. In our attack, we set by default all unrecovered weights to zero, hoping they fall into this category. Clearly it is not the case or the coverage would always sit at 100%. To properly evaluate the quality of the extraction, we must determine whether an unrecovered weight is truly dead. We also aim to understand why some weights are more difficult to recover than others. To this end, we introduce a taxonomy of unrecovered weights that serves both purposes. We classify them into four categories: always-off weights, unreachable inactive weights, unreachable active weights, and query-intensive weights. The first two will turn out to be dead weights. As a preliminary step, Figure 10 illustrates this classification. We then introduce in Section 5.3 a more rigorous set of tests to classify unrecovered weights.
Figure 10: Neuron activation in Model II based on $200,000$ random input points.
Always-off weights. Always-off neurons are clearly visible in the heat map (Figure 10): neurons $\eta^{(2)}_{8}$, $\eta^{(4)}_{5}$, $\eta^{(5)}_{2}$, $\eta^{(5)}_{8}$, $\eta^{(6)}_{4}$, $\eta^{(8)}_{3}$, and $\eta^{(8)}_{6}$ are never activated by any input. As a result, all their associated weights are considered always-off, as well as the weights of neurons in the subsequent layer that connect to them. We can safely classify all of these as dead weights. Similarly, a neuron that is always-on, with an activation of $1.00$, can also be considered dead (see Section 3.3).
Query-intensive weights. The unrecovered space caused by neurons that are almost always-off or almost always-on (respectively meaning that they are inactive/active on almost all the input space) can be reduced by increasing the number of queries. Making more queries allows us to gather more critical points for these neurons, filling in the unrecovered weights that are hard to find. We refer to them as query-intensive weights. However, it is hard to estimate the number of additional queries we need. The number of critical points of a neuron we gather is not clearly related to its activation. This phenomenon is illustrated in Table 2. Sometimes there is a linear relation (layer 2, 6, 8), and sometimes there is no linear relation (layer 4, 7). The correlation indicated is between a layer’s neuron’s $\min(\text{input activation},1-\text{input activation})$, where input activation is the neuron’s number in Table 10, and the ratio of the neuron’s critical points over the critical points of the layer. The former represents how close the neuron is to being always-off or always-on. The closer this number is to $0.5$, the more its extraction is important. The latter shows how hard it is to find critical points for that neuron (a lower ratio denotes higher difficulty in finding critical points for the neuron). To give a concrete example, $\eta^{(7)}_{5}$ and $\eta^{(7)}_{2}$ in Model II have almost the same activation, but we are roughly $2.4$ times more likely to find critical points for the latter than for the former. For this reason, it is hard to predict how much time the attack needs to recover more query-intensive weights. We chose $3,000$ critical points as a middle ground between runtime and coverage.
[…]
Layer 2 3 4 5 6 7 8
Correlation 0.95 0.83 0.58 0.81 0.90 0.65 0.93
Unreachable weights. Suppose that all critical points of a neuron $\eta_{j}^{(i)}$ lie on the inactive side of a neuron $\eta_{k}^{(i-1)}$. In this case, the contribution of $\eta_{k}^{(i-1)}$ to the output of $\eta_{j}^{(i)}$ is always zero during extraction, leaving us with no information about the corresponding weight $a_{j,k}^{(i)}$. As this weight cannot be recovered using the current extraction method, we refer to such weights as unreachable. Unreachable weights arise when a hyperplane from layer $i$ does not intersect with a hyperplane from layer $i-1$, partitioning the input space into three polytopes instead of the usual four. There are two scenarios depending on the sign of $\eta_{j}^{(i)}$. If no region exists where both neurons are active (as illustrated in Figure 12), then $a_{j,k}^{(i)}$ is never activated and can be safely treated as a dead weight. However, if a region does exist where both neurons are active (depicted in grey in Figure 12), then $a_{j,k}^{(i)}$ does contribute and should ideally be recovered. Unlike query-intensive weights, unreachable weights can significantly reduce extraction coverage, as they are not necessarily associated with almost always-off neurons. Nonetheless, almost always-off neurons are more likely to cause unreachable weights (see Table 4). Fortunately, such cases become increasingly rare with higher input dimensions, since the probability that two $(d_{0}-1)$-dimensional hyperplanes do not intersect decreases. For example, a very active unreachable weight, $a^{(7)}_{5,5}$ in Model II, causes a drop in layer coverage to 76.32%, while all other layers maintained values above 88%.
$\eta_{j}^{(i)}$$\eta_{k}^{(i-1)}$++
Figure 11: $a_{j,k}^{(i)}$ is inactive.
$\eta_{j}^{(i)}$$\eta_{k}^{(i-1)}$++
Figure 12: $a_{j,k}^{(i)}$ is active.
5.3 Application to Our Extractions
While helpful, the heat map is not enough to assess the extraction. For example, it cannot tell us about unreachable weights. We use a combination of three activation tests: input activation, “plus-plus” activation, and “plus-minus” activation. Table 3 indicates how those tests yield the taxonomy. We then apply this classification to analyse our extractions.
Input activation. Input activation corresponds to the heat map from Figure 10. We take $200,000$ random input points, and we check for always-off neurons. Since this test is about neurons and not weights, we also need to set the input activation of weights connected to always-off neurons to 00. This test makes a distinction between always-off weights and unreachable inactive weights.
Plus-plus activation. A weight $a_{j,k}^{(i)}$ represents the contribution of $\eta_{k}^{(i-1)}$ to $\eta_{j}^{(i)}$. $a_{j,k}^{(i)}$ contributes to the network when both neurons are active. The plus-plus activation test uses $200,000$ input points to compute the proportion that activates both neurons. This test distinguishes between dead weights and missed weights, and is therefore used to compute coverage.
Plus-minus activation. If $a_{j,k}^{(i)}$ is an active unreachable weight, then there exists no polytope where $\eta_{j}^{(i)}$ is inactive while $\eta_{k}^{(i-1)}$ is active. To test for the existence of such a polytope, we again use $200,000$ random input points. Weights that are always off due to an always-off neuron in the previous layer exhibit a plus-minus activation of 00.
Table 3: Taxonomy of unrecovered weights based on the three tests (Input, (+,+) and (+,-) activations) with some examples.
Taxonomy Conditions
[…]
Dead Weights always-off weights 00 00 $\geq 0$
unreachable inactive weights $>0$ 00 $>0$
Missed Weights unreachable active weights $>0$ $>0$ 00
query-intensive weights $>0$ $>0$ $>0$
Examples
Model II $a_{4,5}^{(7)}$ 00 00 00
$a_{3,6}^{(7)}$ $0.11$ 00 $0.40$
$a_{5,5}^{(7)}$ $0.83$ $0.11$ 00
$a_{2,6}^{(7)}$ $0.11$ $0.11$ $0.48$
We apply this classification to our extractions, showing the results in Table 4. We find that the large majority of unrecovered weights are dead weights: 81%, 83% and 84% for Model I, II and III respectively. This shows that metrics such as the weight recovery rate or the number of neurons fully recovered are not appropriate. Our taxonomy gives insights into coverage values and what we might expect to see if we increase the number of queries. The high coverage of Model I can be explained by the absence of any major missed weights. More queries should help the extraction of Model III reach a high coverage. However, the presence of three major unreachable weights indicates that more queries might not help tremendously the extraction of Model II. To test this we ran the extraction with twice the amount of critical points and observed that indeed Model II’s fidelity barely changed from 74.12% to 74.78% while Model III’s changed more substantially from 71.35% to 79.53%, even though for both models we recovered 0.9% points more weights (see Appendix 0.E for details).
Table 4: Proportion and number of each type of unrecovered weights in all extracted models.
[…]
I always-off weights 00 $72$ $59.50\%$
unreachable inactive weights 00 $26$ $21.49\%$
unreachable active weights $\leq 0.06$ $5$ $4.13\%$
$>0.06$ 00 $0\%$
query-intensive weights $\leq 0.06$ $18$ $14.88\%$
$>0.06$ 00 $0\%$
[…]
unreachable inactive weights 00 $12$ $10.43\%$
unreachable active weights $\leq 0.06$ $5$ $4.35\%$
$>0.06$ $3$ $2.61\%$
query-intensive weights $\leq 0.06$ $7$ $6.09\%$
$>0.06$ $4$ $3.48\%$
[…]
unreachable inactive weights 00 $50$ $32.26\%$
unreachable active weights $\leq 0.06$ 00 $0\%$
$>0.06$ 00 $0\%$
query-intensive weights $\leq 0.06$ $22$ $14.19\%$
$>0.06$ $3$ $1.94\%$
6 Conclusion
In this work, we carried out an in-depth analysis of the approach introduced in [8] for extracting the weights of neural networks, which has since been reused in several follow-up works. We identified critical limitations that prevented this method from successfully recovering the weights of layers beyond the third. Most importantly, we proposed algorithmic improvements for each of these challenges, permitting, for the first time, the extraction of over 99% of the weights of 8 hidden layers networks when each layer is attacked independently. For a $13072-8^{(8)}-$ 1 network, our extracted network’s outputs are within 5% of the target network’s for more than 91% of possible inputs.
Our results establish a new benchmark for signature extraction and open the way to practical attacks on deeper architectures.
Limitations. Going significantly deeper than eight hidden layers is still a challenge, as are floating-point imprecisions in the context of an end-to-end attack. Addressing these limitations, along with a more thorough investigation of how changes in the architecture affect extraction quality, are promising future areas of research.
Appendix
Appendix 0.A Example of a Small Network
We illustrate here the definitions introduced in Section 3.1. Consider the $2-3-2-1$ neural network given by the following layers $A^{(i)}$, $b^{(i)}$:
$A^{(2)}=\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2}&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-4}&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-1}\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-3}&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}4}&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}5}\\ \end{pmatrix},b^{(2)}=\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-2}\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-3}\\ \end{pmatrix}$
@@ Appendix 0.A Example of a Small Network
The architecture of this neural network is depicted in Figure 13. Let the input to the neural network be $v=(x,y)$. The neurons in the first layer output $(\eta^{(1)}_{1}(v),\eta^{(2)}_{1}(v),\eta^{(3)}_{1}(v))=(x+y+1,x-y-2,0.5x+2y-5)$. Depending on $(x,y)$, the ReLU activation function $\sigma$ might set some entries to 00 before passing them to the second layer. For example, when the input is $v=(2,2)$, we have: $F^{(1)}(v)=(5,-2,0)$, $\sigma\circ F^{(1)}(v)=(5,0,0)$, $F^{(2)}(v)=5*(2,-3)+0*(-4,4)+0*(-1,5)+(-2,-3)=(8,-18)$, $\sigma\circ F^{(2)}(v)=(8,0)$, and finally $f(v)=F^{(3)}(v)=8*1+0*(-1)+3=11$. Since $F^{(1)}(v)=(5,-2,0)$, it follows that $v$ is a critical point of $\eta_{3}^{(1)}$.
This network can also be represented as a collection of polytopes (see Figure 13). As can be seen in this figure, critical hyperplanes from the first layer make three straight lines, represented as dotted grey lines: $\eta^{(1)}_{1}:x+y+1$, $\eta^{(1)}_{2}:x-y-2$, and $\eta^{(1)}_{3}:0.5x+2y-5$. The hyperplanes of $\eta_{1}^{(2)}$ and $\eta_{2}^{(2)}$ are shown in their respective colors, with their active side marked by $+$ and their inactive side by --. They indeed break on the grey lines as critical hyperplanes on deeper layers break on all hyperplanes of previous layers: the input to $\eta_{k}^{(i)}$ is $f^{(i-1)}(v)$ rather than $v$ for $i=2,\cdots,r$.
${x}_{1}$${x}_{2}$${\eta}^{(1)}_{1}$${\eta}^{(1)}_{2}$${\eta}^{(1)}_{3}$${\eta}^{(2)}_{1}$${\eta}^{(2)}_{2}$$y$2-3-44-15Input LayerHidden LayersOutput Layer
Figure 13: Network representation of the neural network from Appendix 0.A, with the details of layer $2$ highlighted in colour.
@@
Figure 14: The neural network from Appendix 0.A partitions the input space into 18 distinct polytopes.
@@ 0.C.1 Proof of Lemma 2
Let $x$ be a critical point of layer $i\in\{1,\ldots,r\}$ for neuron $\eta_{k}^{(i)}$ such that $x$ is not a critical point of any other neuron, and $\Delta\in\mathbb{R}^{d_{0}}$. Then, denoting $\Omega_{x}^{(i+1)}=(c_{1},\ldots,c_{d_{i}})$, we have:
$\partial^{2}_{\Delta}f(x)=c_{k}\big{|}\Gamma_{x}^{(i-1)}\Delta\cdot A_{k}^{(i)}\big{|}$
@@
We take $\epsilon$ sufficiently small so that $x+\epsilon\Delta$ and $x-\epsilon\Delta$ are in the neighbourhood of $x$. Thus $x+\epsilon\Delta$ and $x-\epsilon\Delta$ have the exact same activation pattern except for $\eta_{k}^{(i)}$.
$\partial^{2}_{\Delta}f(x)$ $=\frac{1}{\epsilon}(f(x+\epsilon\Delta)+f(x-\epsilon\Delta)-2f(x))$
$=\frac{1}{\epsilon}\bigg{(}G^{(i+1)}\circ\sigma(A^{(i)}F^{(i-1)}(x+\epsilon\Delta)+b^{(i)})$
$+G^{(i+1)}\circ\sigma(A^{(i)}F^{(i-1)}(x-\epsilon\Delta)+b^{(i)})$
@@
$=\frac{1}{\epsilon}\Omega_{x}^{(i+1)}\bigg{(}\sigma(A^{(i)}F^{(i-1)}(x+\epsilon\Delta)+b^{(i)})$
$+\sigma(A^{(i)}F^{(i-1)}(x-\epsilon\Delta)+b^{(i)})$
@@
Let us analyze $C=\sigma(A^{(i)}F_{x}^{(i-1)}(x+\epsilon\Delta)+b^{(i)})+\sigma(A^{(i)}F_{x}^{(i-1)}(x-\epsilon\Delta)+b^{(i)})-2\sigma(A^{(i)}F_{x}^{(i-1)}(x)+b^{(i)})\in\mathbb{R}^{d_{i}\times 1}$.
The coefficients $j\neq k$ of $C$ are zero. Indeed, since $x$, $x+\epsilon\Delta$, and $x-\epsilon\Delta$ belong to the same activation region except for $\eta_{k}^{(i)}$, the computation becomes affine, leading to zero. More precisely:
$C[j]$ $\,=\,[\Gamma_{x}^{(i-1)}(x+\epsilon\Delta)+\gamma_{x}^{(i-1)}]\cdot A_{j}^{(i)}$
$+[\Gamma_{x}^{(i-1)}(x-\epsilon\Delta)+\gamma_{x}^{(i-1)}]\cdot A_{j}^{(i)}$
$-2[\Gamma_{x}^{(i-1)}x+\gamma_{x}^{(i-1)}]\cdot A_{j}^{(i)}$
$=\Gamma_{x}^{(i-1)}x\cdot A_{j}^{(i)}+\epsilon\Gamma_{x}^{(i-1)}\Delta\cdot A_{j}^{(i)}$
$+\,\Gamma_{x}^{(i-1)}x\cdot A_{j}^{(i)}\,-\,\epsilon\Gamma_{x}^{(i-1)}\Delta\cdot A_{j}^{(i)}-2\Gamma_{x}^{(i-1)}x\cdot A_{j}^{(i)}$
$=0\quad\text{when $\eta_{j}^{(i)}$ is active at $x$}$
@@
$F^{(i-1)}(x)\cdot A_{k}^{(i)}\,+\,b^{(i)}_{k}=[\Gamma_{x}^{(i-1)}x+\gamma_{x}^{(i-1)}]\cdot A_{k}^{(i)}\,+\,b^{(i)}_{k}=0$
and since $\Gamma_{x}^{(i-1)}\Delta\cdot A_{k}^{(i)}$ is either positive or negative, we get:
$C[k]=\big{|}\Gamma_{x}^{(i-1)}\Delta\cdot A_{k}^{(i)}\big{|}$
Thus, by multiplying $\Omega^{(i+1)}_{x}$ by $C$, we obtain the expected result.

### v2

@@ Appendix 0.A Example of a Small Neural Network
The architecture of this neural network is depicted in Fig. 10. Let the input to the neural network be $v=(x,y)$. The neurons in the first layer output $(\eta^{(1)}_{1}(v),\eta^{(2)}_{1}(v),\eta^{(3)}_{1}(v))=(x+y+1,x-y-2,0.5x+2y-5)$. Depending on $(x,y)$, the ReLU activation function $\sigma$ might set some entries to 0 before passing them to the second layer. For example, when the input is $v=(2,2)$, we have: $F^{(1)}(v)=(5,-2,0)$, $\sigma\circ F^{(1)}(v)=(5,0,0)$, $F^{(2)}(v)=5*(2,-3)+0*(-4,4)+0*(-1,5)+(-2,-3)=(8,-18)$, $\sigma\circ F^{(2)}(v)=(8,0)$, and finally $f(v)=F^{(3)}(v)=8*1+0*(-1)+3=11$. Since $F^{(1)}(v)=(5,-2,0)$, it follows that $v$ is a critical point of $\eta_{3}^{(1)}$.
This network can also be represented as a collection of polytopes (see Fig. 11). As can be seen in this figure, critical hyperplanes from the first layer make three straight lines, represented as dotted grey lines: $\eta^{(1)}_{1}:x+y+1$, $\eta^{(1)}_{2}:x-y-2$, and $\eta^{(1)}_{3}:0.5x+2y-5$. The hyperplanes of $\eta_{1}^{(2)}$ and $\eta_{2}^{(2)}$ are shown in their respective colors, with their active side marked by ++ and their inactive side by -. They indeed break on the grey lines as critical hyperplanes on deeper layers break on all hyperplanes of previous layers: the input to $\eta_{k}^{(i)}$ is $F^{(i-1)}(v)$ rather than $v$ for $i=2,\cdots,r$.
${x}_{1}$${x}_{2}$${\eta}^{(1)}_{1}$${\eta}^{(1)}_{2}$${\eta}^{(1)}_{3}$${\eta}^{(2)}_{1}$${\eta}^{(2)}_{2}$yy2-3-44-15Input LayerHidden LayersOutput Layer
Figure 10: Network representation of the neural network example, with the details of layer 22 highlighted in colour.
@@
Figure 11: The neural network example partitions the input space into 18 distinct polytopes.
@@ 0.C.1 Proof of Lemma 2
Let $x$ be a critical point of layer $i\in\{1,\ldots,r\}$ for neuron $\eta_{k}^{(i)}$ such that $x$ is not a critical point of any other neuron, and $\Delta\in\mathbb{R}^{d_{0}}$. We have that:
$\partial^{2}_{\Delta}f(x)=c_{k}^{(i)}\big|A_{k}^{(i)}\cdot\Gamma_{x}^{(i-1)}\Delta\big|$
@@
We take $\epsilon_{2}$ sufficiently small so that $x+\epsilon_{2}\Delta$ and $x-\epsilon_{2}\Delta$ are in the neighbourhood of $x$. Thus $x+\epsilon_{2}\Delta$ and $x-\epsilon_{2}\Delta$ have the exact same activation pattern except for $\eta_{k}^{(i)}$. Let’s write $G^{(i)}$ for the composition of layers starting from $i$, i.e. $f=G^{(i+1)}\circ F^{(i)}$.
$\partial^{2}_{\Delta}f(x)$ $=\frac{1}{\epsilon_{1}}(f(x+\epsilon_{2}\Delta)+f(x-\epsilon_{2}\Delta)-2f(x))$
$=\frac{1}{\epsilon_{1}}\bigg(G^{(i+1)}\circ\sigma(A^{(i)}F^{(i-1)}(x+\epsilon_{2}\Delta)+b^{(i)})$
$+G^{(i+1)}\circ\sigma(A^{(i)}F^{(i-1)}(x-\epsilon_{2}\Delta)+b^{(i)})$
@@
$=\Omega_{x}^{(i)}\bigg(\sigma(A^{(i)}F^{(i-1)}(x+\epsilon_{2}\Delta)+b^{(i)})$
$+\sigma(A^{(i)}F^{(i-1)}(x-\epsilon_{2}\Delta)+b^{(i)})$
@@
Where $\Omega_{x}^{(i)}$ is $\frac{1}{\epsilon_{1}}(c_{1}^{(i)},\ldots,c_{d_{i}}^{(i)})$ for some values $c_{n}\in\mathbb{R}$. Let us analyze $C=\sigma(A^{(i)}F_{x}^{(i-1)}(x+\epsilon_{2}\Delta)+b^{(i)})+\sigma(A^{(i)}F_{x}^{(i-1)}(x-\epsilon_{2}\Delta)+b^{(i)})-2\sigma(A^{(i)}F_{x}^{(i-1)}(x)+b^{(i)})\in\mathbb{R}^{d_{i}\times 1}$.
The coefficients $j\neq k$ of $C$ are zero. Indeed, since $x$, $x+\epsilon_{2}\Delta$, and $x-\epsilon_{2}\Delta$ belong to the same activation region except for $\eta_{k}^{(i)}$, the computation becomes affine, leading to zero. More precisely:
$C[j]$ $\,=\,A_{j}^{(i)}\cdot[\Gamma_{x}^{(i-1)}(x+\epsilon_{2}\Delta)+\gamma_{x}^{(i-1)}]$
$+A_{j}^{(i)}\cdot[\Gamma_{x}^{(i-1)}(x-\epsilon_{2}\Delta)+\gamma_{x}^{(i-1)}]$
$-A_{j}^{(i)}\cdot 2[\Gamma_{x}^{(i-1)}x+\gamma_{x}^{(i-1)}]$
$=A_{j}^{(i)}\cdot\Gamma_{x}^{(i-1)}x+A_{j}^{(i)}\cdot\epsilon_{2}\Gamma_{x}^{(i-1)}\Delta$
$+\,A_{j}^{(i)}\cdot\Gamma_{x}^{(i-1)}x\,-\,A_{j}^{(i)}\cdot\epsilon_{2}\Gamma_{x}^{(i-1)}\Delta-A_{j}^{(i)}\cdot 2\Gamma_{x}^{(i-1)}x$
$=0\quad\text{when $\eta_{j}^{(i)}$ is active at $x$.}$
@@
$A_{k}^{(i)}\cdot F^{(i-1)}(x)\,+\,b^{(i)}_{k}=A_{k}^{(i)}\cdot[\Gamma_{x}^{(i-1)}x+\gamma_{x}^{(i-1)}]\,+\,b^{(i)}_{k}=0$
and since $A_{k}^{(i)}\cdot\Gamma_{x}^{(i-1)}\Delta$ is either positive or negative, we get:
$C[k]=\big|A_{k}^{(i)}\cdot\Gamma_{x}^{(i-1)}\Delta\big|$
Thus, by multiplying $\Omega^{(i)}_{x}$ by $C$, we obtain the expected result.
Intuition behind the above proof is given in Fig. 12 below.
[…]
(a) $\eta(x)\geq 0$ for $x\geq x^{*}$
[…]
(b) $\eta(x)\geq 0$ for $x\leq x^{*}$
Figure 12: Illustration of how a neuron’s activation change at a critical point creates a detectable break in the network’s output, allowing the extraction of the neuron. The solid line shows the true output $f(x)$, while the dotted line shows the hypothetical output if the neuron $\eta$ were always inactive. At the critical point $x^{*}$, $\eta$ flips its state, creating a change in slope. Two cases can occur: (a) the neuron becomes active ($\eta(x)\geq 0$ for $x\geq x^{*}$), in which case the solid line bends away from the dotted line; or (b) the neuron becomes inactive ($\eta(x)\geq 0$ for $x\leq x^{*}$), in which case the solid line aligns with the dotted line after $x^{*}$. The slopes before and after $x^{*}$ (red arrows) differ by the absolute weight of $\eta$, shown by the blue arrow. This indicates how much the neuron contributes, hence its weights, although an additional step is required to determine the correct sign of the contribution.